OpenStudy (anonymous):

Let u = <3, -1>, v = <-6, -6>. Find 9u + 2v.

Parth (parthkohli):

Can you first find \(9\bf \vec u \)?

OpenStudy (anonymous):

@ParthKohli would it be 9(3+-6)=-27?

OpenStudy (anonymous):

if so 2v =-14?

Parth (parthkohli):

Erm, well no...

OpenStudy (anonymous):

dont add them together, leave it in the form of a vector

Parth (parthkohli):

\[9{\bf \vec u} = 9 \times \langle 3,-1) = \langle 9\times 3, 9 \times -1\rangle = \langle 27,-9\rangle \]

Parth (parthkohli):

Can you do it for \(2\bf \vec v\)? We'll add them together afterwards.

OpenStudy (anonymous):

2v= 2(-6,-6)= (2x-6, 2 x -6)= (-12,-12)

Parth (parthkohli):

Nailed it! Now add them.

OpenStudy (anonymous):

15,-21

Parth (parthkohli):

Whoa, that's correct :-D

OpenStudy (anonymous):

Thanks for your help

Parth (parthkohli):

No problem!